# Fantastic Weights and How to Find Them: Where to Prune in Dynamic Sparse Training

### v1

@@ 5.4 Structural Similarity
Finally, we investigate whether the studied criteria are diverse in terms of the selected weights for removal. We fix the dataset to CIFAR10 and analyze the moment in which the sparse connectivity is updated for the first time. At such a point, given the same sparse initialization and random seed, all the pruning criteria share the same network state. Consequently, we can analyze the difference between the sets selected for removal by each criterion. We do this by computing their Jaccard index (Equation 1) and present the averaged results of 5 different runs in Figure 5(a) (See also Appendix I).
@@ 6 Conclusion
Limitations & Future Work. Our research was conducted mainly on datasets from computer vision and one tabular dataset. It would be interesting to verify how our findings translate to large language models. The computed structural insights are based on set similarity and disregard the information about the value of the parameters. Additional topographic insights could be provided to incorporate this information and analyze the graph structure of the found connectivity. Finally, much work still needs to be done on the hardware side to truly speed up training times and decrease memory requirements (see Appendix E). We do not see any direct social or ethical broader impact of our work.
@@ A.1 Model Architectures
For the small-MLP setup, we use an MLP with 2 hidden layers with dimension sizes of $256$. For the large MLP, we increase the number of neurons in the first hidden layer to $1024$, and to $512$ in the second. The small-CNN is a standard ConvNet architecture consisting of three convolutional layers with kernel size $3$, stride $1$, and padding $1$, inspired by the standard ConvNet presented in section 8.2.3 of [7]. The spatial dimension reduction is implemented by MaxPooling with size $2$. In all cases, we use ReLU activations. The architectures are summed up in Table 2. The ResNet-56 and ResNet-50 models follow their standard specifications [21]. The EfficientNet is the EfficientNet-B0 model [52] (we use the implementation from pytorch models), and follows the standard architecture. The VGG-16 model is modified to be used with CIFAR-10 and uses BatchNorm. The implementation of VGG-16, as well as the LeNet-5-Caffe model is adapted from [33].
@@ A.2 Training Regime
DST Settings specific for dynamic sparse training: We use the ER initialization for the distribution of sparsity levels in MLP, and the ERK initialization for the convolutional models. We start with an initial pruning fraction $\rho$ of $0.5$, which is decayed during training using a cosine decay schedule (see the main text). The default topology update period $\Delta t$ is set to $800$, but in the experiments for Q2 we vary this hyperparameter. Note that we also keep the update period $\Delta t=800$ for the ImageNet, since together with batch size $256$ it matches the number of training examples seen by the model before the next topology update reported in [12].
@@
We perform $5$ runs for each setup, except for the ImageNet dataset, in which we use $3$ runs. We do not fine-tune anything specific for any model. In all the experiments performed, at the beginning of the code, we fixed all the random seeds (random.seed, numpy.random, torch.manual_seed, torch.cuda.manual_seed), and set the cudnn backend to deterministic. The DST pruning criteria (if not used with random growth) are deterministic and for the same seed start with the same sparse initialization. While computing the Jaccard Index between the end masks we use only one worker for data-loading. In terms of reproducibility, we provide the environment setup but also perform our experiments in a Singularity container, for which we provide the definition file in the repository. We provide the hyperparameter setups for all the main-text experiments in the code  [^9] Under the specs directory in the code: https://github.com/alooow/fantastic_weights_paper/tree/main/specs.
@@
Our code is based on the publicly available repositories for [33] and [11] available at https://github.com/VITA-Group/Random_Pruning and https://github.com/TimDettmers/sparse_learning, respectively. The used sparse-learning library (of [11]) is published under the MIT licence.  [^10] We include the copy of that licence in the sparse-learning directory in our code.
@@
ImageNet image 1,431,167 3x469x387131313This is the average resolution. Images are randomly cropped to size 3x224x224, as is standard practice. 1000 90%-3%-7% [^14] Percentages are rounded. We use the standard train/validation/test split. We do not use the test set and evaluate on the validation set. [47] image-net.org/challenges/LSVRC/2012/
@@
Tiny-ImageNet image 100000 3x64x64 200 80%-10%-10% [^16] We use the standard train/validation/test split. We use the validation set as the test set (we did not optimize any hyperparameters using the validation set).  [28] cs231n.stanford.edu/tiny-imagenet-200.zip
@@ Appendix B Sensitivity and Reciprocal Sensitivity
Appendix C Additional Performance Plots
@@
We observe that again the differences between the studied methods seem to be more visible in the low-density regime. In that setup, the $\mathcal{C}_{SNIP}$ criterion is achieving worse test accuracy than the $\mathcal{C}_{MEST}$ and the magnitude-based methods. The $\mathcal{C}_{RSensitivity}$ is also perform poorly on the CNN-CIFAR10 model. We can also see that the behaviour for VGG-16 network observed in the plots in the main text is not due to the change of dataset, but rather the change of the model (since ResNet56 on CIFAR100 has similar results to ResNet56 on CIFAR10 from the main text).
@@
Appendix D In-Time-Over-Parametrization (ITOP) Study
@@
We study the relationship between ITOP, update period and the pruning criteria on the example of the large-MLP model on CIFAR10 with density $0.1$. The results are shown in Figure 10. There are some interesting insights gained by this analysis.
Figure 10: The update period $\Delta t$ versus the ITOP ratio obtained at the end of the training for different pruning criteria on the large-MLP. The left plot corresponds to random growth, the right plot corresponds to gradient growth.
First, we observe that when paired with random growth, the choice of pruning criteria leads to almost identical final ITOP ratios. (Note that in the left plot the points for different pruning criteria overlap, yet, we do verify that they are not exactly equal). This is expected, because in such a case the growth criterion is independent of the chosen set of weights to prune. Hence the increase in ITOP is only due to growth criterion, which (being random selection) is not affected by the pruning criterion. This is in contrast to gradient growth, in which the change in the pruned weights can impose a change to the gradient flow in the model and hence impact the rate of the exploration. Aside from this observation we confirm that gradient growth has in general a lower ITOP ratio than random (which was shown in  [37]).
Secondly, in the gradient growth setup, the magnitude criteria, as well as $\mathcal{C}_{MEST}$, have a much higher ITOP ratio for small update periods than the gradient-based pruning criteria. This indicates that they perform more exploration, which perhaps may be the reason behind their better performance observed in Figure 4 in the main text.
Finally, we notice that final ITOP ratio close to 1 is not necessary for good generalization. Observe that for $\Delta t=800$ we get the best validation accuracy for random growth, MLP, CIFAR10 (Figure 5 in our paper), but it only reaches an ITOP ratio of $0.7$. Even with a final ITOP ratio of just $0.2$ (for $\Delta t=6400$) sparse models can beat the performance of dense models.
Appendix E Sparse Hardware and Software Support
@@
Appendix F Gradient Based Methods and Batch Size
@@
In Section 5.2 in the main text we observe that in general, the magnitude-based pruning criteria perform better in sparse regimes than the gradient-based ones. We loosely hypothesise that this may be due to high variance in the gradient. The gradient is computed for a batch (using SGD). Hence it might happen in Taylor-approximation based pruning (e.g. SNIP) that it is not a good indicator of the change in the loss before and after the pruning. At the same time computing the gradient over the entire training set is not feasible in DST, since it would significantly increase the time complexity of each mask update. To investigate this intuition we perform a small study on the LeNet-5-Caffe model (trained on FashionMNIST), in which we increase the batch size almost $10$ times up to $1024$. We compare the obtained results with the standard setup with batch size $128$. We also decrease the update period from $800$ to $100$, so that in both setups the DST updates are made after seeing the same number of training examples. We report the results in Figure 11.
Interestingly, we observe that for batch size $1024$ $\mathcal{C}_{SNIP}$ indeed seems to perform better. $\mathcal{C}_{RSensitivity}$, on the other hand, becomes significantly unstable, and $\mathcal{C}_{MEST}$ is still indistinguishable from magnitude pruning. Moreover, the general performance with batch $1024$ is worse than with batch $128$ (this is often the case in deep learning, where the stochastic nature of SGD helps with generalization). Those results may indicate that there is indeed a relationship between the noisiness of the gradients and the DST performance. However, this hypothesis needs to be further investigated in order to be evaluated with certainty.
@@
Figure 11: The performance of different pruning criteria on LeNet-5-Caffe on FashionMNIST. The first row corresponds to random growth, the second row corresponds to gradient growth. The left column indicates batch of size $128$, the right column indicates batch of size $1024$.
Appendix G Fine-Tuning EfficientNet and the Choice of Pruning Criterion
@@
In all experiments in the main text we perform the training from scratch (no pretraining is applied). This includes the EfficientNet on the Tiny-ImageNet from the main text. In this section we also analyze the difference in the performance of the pruning criteria when evaluated in a fine-tuning scenario. In such a setup, we initialize the weights of the EfficientNet to use the pretrained parameters from ImageNet. We only re-initialize the linear projection in the head of the model (which needs to have the output dimension changed from 1000 to 200, since there are 200 classes in Tiny-ImageNet). We keep all hyperparametrs the same as in the training-from-scratch setup (see Appendix A). We compare the results in Figure 12
@@
We observe that in the setup that uses the pretrained weights the $\mathcal{C_{SNIP}}$ criterion performs much better than when used training from scratch. We hypothesize that perhaps the use of the Taylor approximation of the change in the Loss function (used in the SNIP criterion) allows to not diverge far from the already quite well-established solution from the pretraining. The purely magnitude-based methods, on the other hand, disregard the potential change in the loss they may cause, and hence do not successfully leverage the information from pretraining. However, at the same time, the general performance of all DST methods when starting with pretrained parameters on ImageNet is much worse in the low-density regime, in comparison to their DST counterparts trained from scratch. This may suggest the the DST framework in the fine-tuning scenario is heavily biased by the pretrained weights and potentially does not efficiently explore possible connectivity patterns. This results strongly suggest that one should be careful not to transfer the findings on the DST training to fine-tuning scenarios. In the second case, more research is needed to understand the behaviour of the DST algorithms.
@@
Figure 12: The performance of different pruning criteria on the EfficientNet architecture trained from scratch (left) and fine-tuned on the weights from training on ImageNet (right). The first row corresponds to random growth, the second row corresponds to gradient growth. We can observe that in the setup that uses the pretrained weights (right column), the $\mathcal{C_{SNIP}}$ criterion performs much better than in training from scratch. However, at the same time, the general performance of all DST methods is much worse in the low-density regime, in comparison to the training from scratch.
Appendix H Global versus Local Pruning

### v2

@@ 5.4 Structural Similarity
Finally, we investigate whether the studied criteria are diverse in terms of the selected weights for removal. We fix the dataset to CIFAR10 and analyze the moment in which the sparse connectivity is updated for the first time. At such a point, given the same sparse initialization and random seed, all the pruning criteria share the same network state. Consequently, we can analyze the difference between the sets selected for removal by each criterion. We do this by computing their Jaccard index (Equation 1) and present the averaged results of 5 different runs in Figure 5(a) (See also Appendix K standard deviations and more plots).
@@ 6 Conclusion
Limitations & Future Work. Our research was conducted mainly on datasets from computer vision and one tabular dataset. It would be interesting to verify how our findings translate to large language models. The computed structural insights are based on set similarity and disregard the information about the value of the parameters. Additional topographic insights could be provided to incorporate this information and analyze the graph structure of the found connectivity. Finally, much work still needs to be done on the hardware side to truly speed up training times and decrease memory requirements (see Appendix F). We do not see any direct social or ethical broader impact of our work.
Acknowledgements
The research of Aleksandra Nowak and Jacek Tabor has been supported by the flagship project entitled Artificial Intelligence Computing Center Core Facility from the DigiWorld Priority Research Area under the Strategic Programme Excellence Initiative at Jagiellonian University. The work of J. Tabor was supported by the National Centre of Science (Poland) Grant No. 2021/41/B/ST6/01370. The research of Bram Grooten was funded by the project AMADeuS (with project number 18489) of the Open Technology Programme, which is partly financed by the Dutch Research Council (NWO). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2022/015887.
@@ A.1 Model Architectures
For the small-MLP setup, we use an MLP with 2 hidden layers with dimension sizes of $256$. For the large MLP, we increase the number of neurons in the first hidden layer to $1024$, and to $512$ in the second. The small-CNN is a standard ConvNet architecture consisting of three convolutional layers with kernel size $3$, stride $1$, and padding $1$, inspired by the standard ConvNet presented in section 8.2.3 of [7]. The spatial dimension reduction is implemented by MaxPooling with size $2$. In all cases, we use ReLU activations. The architectures are summed up in Table 2. The ResNet-56 and ResNet-50 models follow their standard specifications [22]. The EfficientNet is the EfficientNet-B0 model [57] (we use the implementation from pytorch models), and follows the standard architecture. The VGG-16 model is modified to be used with CIFAR-10 and uses BatchNorm. The implementation of VGG-16, as well as the LeNet-5-Caffe model, is adapted from [34].
@@ A.2 Training Regime
DST Settings specific for dynamic sparse training: We use the ER initialization for the distribution of sparsity levels in MLP, and the ERK initialization for the convolutional models. We start with an initial pruning fraction $\rho$ of $0.5$, which is decayed during training using a cosine decay schedule (see the main text). The default topology update period $\Delta t$ is set to $800$, but in the experiments for Q2 we vary this hyperparameter. Note that we also keep the update period $\Delta t=800$ for the ImageNet, since together with batch size $256$ it matches the number of training examples seen by the model before the next topology update reported in [12]. For completeness, we also provide the DST hyperparameter setups used to obtain the Figures in the main text in Table 3.
Table 3: The DST hyperparameters used to obtain the results presented in the main-text experiments.
Experiment Prune Fraction Prune Fraction Decay Scheduler Density Update Period
Fig. 1 0.5 cosine 0.05, 0.06, 0.07, 0.08, 0.09, 800
0.1, 0.15, 0.2, 0.3, 0.4, 0.5
Fig. 2 0.5 cosine 0.2 800
Fig. 4 0.5 cosine 0.2, 0.1 25, 50, 100, 200, 400,
800, 1600,3200,6400
Fig. 5 0.5 cosine 0.5 800
Fig. 6a 0.5 cosine 1.0 800 (measured only once)
Fig. 6b 0.5 cosine 0.2 800
@@
We perform $5$ runs for each setup, except for the ImageNet dataset, in which we use $3$ runs. We do not fine-tune anything specific for any model. In all the experiments performed, at the beginning of the code, we fixed all the random seeds (random.seed, numpy.random, torch.manual_seed, torch.cuda.manual_seed), and set the cudnn backend to deterministic. The DST pruning criteria (if not used with random growth) are deterministic and for the same seed start with the same sparse initialization. While computing the Jaccard Index between the end masks we use only one worker for data-loading. In terms of reproducibility, we provide the environment setup but also perform our experiments in a Singularity container, for which we will provide the definition file in the repository. We provide the hyperparameter setups for all the main-text experiments in the code. [^10] Under the specs directory in the code.
@@
Our code is accessible at https://github.com/alooow/fantastic_weights_paper. It is based on the publicly available repositories for [34] and [11] available at https://github.com/VITA-Group/Random_Pruning and https://github.com/TimDettmers/sparse_learning, respectively. The used sparse-learning library (of [11]) is published under the MIT license. [^11] We include the copy of that licence in the sparse-learning directory in our code.
@@
ImageNet image 1,431,167 3x469x387141414This is the average resolution. Images are randomly cropped to size 3x224x224, as is standard practice. 1000 90%-3%-7% [^15] Percentages are rounded. We use the standard train/validation/test split. We do not use the test set and evaluate on the validation set. [49] image-net.org/challenges/LSVRC/2012/
@@
Tiny-ImageNet image 100000 3x64x64 200 80%-10%-10% [^17] We use the standard train/validation/test split. We use the validation set as the test set (we do not optimize any hyperparameters using the validation set).  [29] cs231n.stanford.edu/tiny-imagenet-200.zip
@@ Appendix B Sensitivity and Reciprocal Sensitivity
Appendix C Results On Text Data
In addition to the tasks studied in the main text, we conduct an evaluation of the pruning criteria on the fine-tuning task of ROBERTa Large [39] with approximately 354 million parameters, utilizing the CommonsenseQA dataset [56] adapted from the Sparsity May Cry (SMC) Benchmark [35].
We use the exact same setup as in the SMC-Benchmark by first performing magnitude pruning for the sparse initialization and then using DST during the fine-tuning phase. We also use the same hyperparameters. Please note that this is a different configuration from the one used in the main text, in which we trained all the models from scratch (in spirit, the CommonsenseQA study is more similar to the EfficientNet fine-tuning experiment from Appendix H). We report the results in Table 5 and Table 6 for the random and gradient growth, respectively.
Similar to the findings of  [35], we observe that in this scenario, the achievable sparsity without a significant performance decline is notably lower than in vision or tabular data scenarios. Concerning random growth, the $\mathcal{C}_{MEST}$ criterion appears to be the most suitable choice. For gradient growth with a density of 0.8, the $\mathcal{C}_{RSensitivity}$ and $\mathcal{C}_{SET}$ criteria initially demonstrate strong performance but are eventually surpassed by the $\mathcal{C}_{magnitude}$ and $\mathcal{C}_{MEST}$ criteria at a density of 0.7. Additionally, we note a significantly higher variance in the outcomes across all criteria compared to vision tasks.
Furthermore, we conducted a brief exploration of update period values for this problem (refer to Figure 9). Notably, we observe that overly frequent updates do not yield beneficial results. Concerning gradient growth, the most effective update period value is consistently around $\Delta t=500$ for nearly all criteria. For random growth higher update period values generally lead to improved outcomes.
At the same time, please note that the studied pruning criteria had been evaluated either on tabular or vision datasets in most of the works that have introduced them or used them (e.g.,  [42, 64, 45, 12, 11]). Large language models typically have a different structure and are based on the attention mechanism. In consequence, the applicability of sparse training in such a setup is still an open area of research [35, 55]. This motivated our choice of fixing our focus on tabular and vision models in the main text.
Table 5: The mean accuracy results for the CommonsenseQA task on ROBERTa Large Model from the SMC-Benchmark (Liu et al. 2023). We use the same hyperparameter setup as in (Liu et al. 2023) and vary only the density and pruning criterion. The results below were obtained for random growth. In brackets, we report the standard deviation from 3 runs. We bold out the best result for each density and underline the second-best one.
random growth
0.9 0.8 0.7 0.6 0.5
$\mathcal{C}_{magnitude}$ 73.80 (0.52) 37.57 (25.25) 26.53 (6.71) 19.80 (3.13) 20.60 (0.69)
$\mathcal{C}_{SET}$ 74.40 (0.60) 49.50 (26.98) 21.00 (3.99) 20.13 (1.95) 18.63 (1.44)
$\mathcal{C}_{MEST}$ 74.77 (0.67) 52.30 (29.79) 26.90 (7.20) 19.77 (0.49) 20.73 (0.96)
$\mathcal{C}_{RSensitivity}$ 25.87 (8.36) 22.13 (1.69) 23.27 (0.85) 21.37 (1.66) 21.20 (0.53)
$\mathcal{C}_{SNIP}$ 56.57 (1.82) 34.20 (9.11) 19.57 (1.31) 21.60 (0.95) 19.17 (1.47)
Table 6: The mean accuracy results for the CommonsenseQA task on ROBERTa Large Model from the SMC-Benchmark (Liu et al. 2023). We use the same hyperparameter setup as in (Liu et al. 2023) and vary only the density and pruning criterion. The results below were obtained for gradient growth. In brackets, we report the standard deviation from 3 runs. We bold out the best result for each density and underline the second-best one.
gradient growth
0.9 0.8 0.7 0.6 0.5
$\mathcal{C}_{magnitude}$ 76.00 (0.46) 55.80 (29.51) 37.37 (18.23) 24.27 (2.94) 19.33 (0.35)
$\mathcal{C}_{SET}$ 76.30 (0.75) 68.43 (6.15) 28.57 (1.85) 22.60 (1.47) 22.17 (0.93)
$\mathcal{C}_{MEST}$ 74.73 (0.15) 53.90 (29.27) 37.53 (8.45) 24.27 (2.21) 19.63 (1.12)
$\mathcal{C}_{RSensitivity}$ 75.83 (0.55) 69.60 (6.70) 25.43 (5.65) 20.67 (2.81) 21.57 (1.82)
$\mathcal{C}_{SNIP}$ 65.50 (2.27) 48.20 (7.79) 32.70 (5.98) 23.17 (4.82) 20.73 (0.61)
Figure 9: The study of the effect of the change of update period $\Delta t$ on the performance in the CommonsenseQA task on ROBERTa Large. Left: random growth, right: gradient growth.
Appendix D Additional Performance Plots
@@
We observe that again the differences between the studied methods seem to be more visible in the low-density regime. In that setup, the $\mathcal{C}_{SNIP}$ criterion is achieving worse test accuracy than the $\mathcal{C}_{MEST}$ and the magnitude-based methods. The $\mathcal{C}_{RSensitivity}$ also performs poorly on the CNN-CIFAR10 model. We can see that the behavior for the VGG-16 network observed in the plots in the main text is not due to the change of dataset, but rather the change of the model (since ResNet56 on CIFAR100 has similar results to ResNet56 on CIFAR10 from the main text).
@@
Appendix E In-Time-Over-Parametrization (ITOP) Study
@@
We study the relationship between ITOP, update period, and the pruning criteria on the example of the large-MLP model on CIFAR10 with density $0.1$. The results are shown in Figure 11. There are some interesting insights gained by this analysis.
Figure 11: The update period $\Delta t$ versus the ITOP ratio obtained at the end of the training for different pruning criteria on the large MLP. The left plot corresponds to random growth, the right plot corresponds to gradient growth.
First, we observe that when paired with random growth, the choice of pruning criteria leads to almost identical final ITOP ratios. (Note that the points for different pruning criteria overlap in the left plot, yet we verify that they are not exactly equal). This is expected because, in such a case, the growth criterion is independent of the chosen set of weights to prune. Hence, the increase in ITOP is only due to the growth criterion, which (being random selection) is not affected by the pruning criterion. This is in contrast to gradient growth, in which the change in the pruned weights can impose a change to the gradient flow in the model and hence impact the rate of the exploration. Aside from this observation, we confirm that gradient growth has, in general, a lower ITOP ratio than random (which was shown in  [38]).
Secondly, in the gradient growth setup, the magnitude criteria and $\mathcal{C}_{MEST}$ have a much higher ITOP ratio for small update periods than the gradient-based pruning criteria. This indicates that they perform more exploration, which perhaps may be the reason behind their better performance observed in Figure 4 in the main text.
Finally, we notice that final ITOP ratio close to $1$ is not necessary for good generalization. Observe that for $\Delta t=800$, we get the best validation accuracy for random growth, MLP, and CIFAR10 (Figure 5 in our paper), but it only reaches an ITOP ratio of $0.7$. Even with a final ITOP ratio of just $0.2$ (for $\Delta t=6400$), sparse models can beat the performance of dense models.
Appendix F Sparse Hardware and Software Support
@@
Appendix G Gradient Based Methods and Batch Size
@@
In Section 5.2 in the main text, we observe that, in general, the magnitude-based pruning criteria perform better in sparse regimes than the gradient-based ones. Note that the gradient is computed for a batch (using SGD), and hence, small batch sizes might introduce more significant errors in Taylor-approximation based pruning (e.g., SNIP). At the same time, computing the gradient over the entire training set is not feasible in DST since it would significantly increase the time complexity of each mask update.
To investigate the impact of the batch size on our results, we perform a small study on the LeNet-5-Caffe model (trained on FashionMNIST), in which we increase the batch size almost $10$ times up to $1024$. At the same time, we increase the learning rate up to $0.08$ from $0.01$ so that the learning-rate-to-batch-size ratio stays approximately the same (as suggested by the scaling laws, for instance, in  [51]). We compare the obtained results with the standard setup with a batch size of $128$. We also decrease the update period from $800$ to $100$ so that in both setups, the DST updates are made after seeing the same number of training examples. We report the results in Figure 12.
Interestingly, we observe that for batch size $1024$ $\mathcal{C}_{SNIP}$ indeed seems to perform slightly better. $\mathcal{C}_{RSensitivity}$ and $\mathcal{C}_{SET}$, on the other hand, become significantly unstable, and $\mathcal{C}_{MEST}$ is still indistinguishable from magnitude pruning. Those results may indicate that the impact of the batch size, although existent, is not significant. Furthermore, as we will see in the next paragraph, increasing only the batch size of the DST update is also not beneficial.
Increasing Only the Batch Size of the DST update.
In addition to the above-conducted experiments, we examine an ablation in which we increase the batch size to $1024$ only when computing the gradient used for the DST update. During normal steps of optimization, we use the standard choice from previous experiments ($128$). The results are in Figure 13. We observe that it does not enhance performance. Let us note, however, that, in general, the impact of the batch size still needs additional investigation in larger models and datasets in order to be evaluated with certainty.
@@
Figure 12: The performance of different pruning criteria on LeNet-5-Caffe on FashionMNIST. The first row corresponds to random growth, the second row corresponds to gradient growth. The left column indicates a batch of size $128$, and the right column indicates a batch of size $1024$. The black dashed line in both plots is the performance of the dense model from the main text.
Figure 13: The performance of MLP-CIFAR10 and LeNet5-Caffe-FashionMNIST from the main text experiments (first and third column) and the performance when using an increased batch only for the DST update (second and last column).
Appendix H Fine-Tuning EfficientNet and the Choice of Pruning Criterion
@@
In all experiments in the main text, we perform the training from scratch (no pretraining is applied). This includes the EfficientNet on the Tiny-ImageNet from the main text. In this section, we also analyze the difference in the performance of the pruning criteria when evaluated in a fine-tuning scenario. In such a setup, we initialize the weights of the EfficientNet to use the pretrained parameters from ImageNet. We only re-initialize the linear projection in the head of the model (which needs to have the output dimension changed from 1000 to 200 since there are 200 classes in Tiny-ImageNet). We keep all hyperparameters the same as in the training-from-scratch setup (see Appendix A). We compare the results in Figure 14
@@
We observe that in the setup that uses the pretrained weights, the $\mathcal{C_{SNIP}}$ criterion performs much better than when training from scratch. We hypothesize that perhaps the use of the Taylor approximation of the change in the Loss function (used in the SNIP criterion) allows not to diverge far from the already quite well-established solution from the pretraining. The purely magnitude-based methods, on the other hand, disregard the potential change in the loss they may cause and hence, do not successfully leverage the information from pretraining. However, at the same time, the general performance of all DST methods when starting with pretrained parameters on ImageNet is much worse in the low-density regime compared to their DST counterparts trained from scratch. This may suggest that the DST framework in the fine-tuning scenario is heavily biased by the pretrained weights and potentially does not efficiently explore possible connectivity patterns. These results suggest that one should be careful not to transfer the findings of the DST training to fine-tuning scenarios. In the second case, more research is needed to understand the behavior of the DST algorithms.
@@
Figure 14: The performance of different pruning criteria on the EfficientNet architecture trained from scratch (left) and fine-tuned on the weights from training on ImageNet (right). The first row corresponds to random growth, and the second row corresponds to gradient growth. We can observe that in the setup that uses the pretrained weights (right column), the $\mathcal{C_{SNIP}}$ criterion performs much better than in training from scratch. However, at the same time, the general performance of all DST methods is much worse in the low-density regime compared to the training from scratch.
Appendix I Global versus Local Pruning